# A Survey on 3D Gaussian Splatting

### v4

@@ 6 Performance Comparison
In this section, we provide more empirical evidence by presenting the performance of several 3D GS algorithms that we previously discussed. The diverse applications of 3D GS across numerous tasks, coupled with the custom-tailored algorithmic designs for each task, render a uniform comparison of all 3D GS algorithms across a single task or dataset impracticable. Therefore, drawing from our analysis in Sec.  5, we have chosen several representative tasks within the 3D GS domain for an in-depth performance evaluation. The performance scores are primarily sourced from the original papers, except where indicated otherwise.
@@ 6.1 Performance Benchmarking: Localization
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent 3D GS based algorithms [107, 106, 105, 104, 113] and six typical SLAM methods [256, 257, 258, 259, 260, 261]
@@
$\bullet$ Result: As shown in Table I, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, HFSLAM [113] achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.25cm compared to the previous state-of-the-art (SOTA) [260]. We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
@@ 6.2 Performance Benchmarking: Static Scenes
$\bullet$ Dataset: The same dataset as in Sec. 6.1, i.e., Replica [255], is used for performance comparison.
@@
$\bullet$ Result: Table II shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM [107] establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM [260], GSSLAM [106] is about 578 times faster in achieving very competitive accuracy. In contrast to previous method [260] that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
@@ 6.3 Performance Benchmarking: Dynamic Scenes
$\bullet$ Dataset: D-NeRF [170] dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in the monocular setting.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers that model dynamic scenes with 3D GS [119, 95, 176, 93], as well as six NeRF based approaches [170, 265, 33, 266, 173, 267].
@@
$\bullet$ Result: From Table III we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS [10] fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS [94] outperforms the SOTA method, FFDNeRF [173], by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
TABLE III: Quantitative rendering results6.3) on D-NeRF [170], in terms of PSNR, SSIM, and LPIPS. denotes numbers taken from [95].
@@
TABLE IV: Quantitative rendering results6.4) on nuScenes [243], in terms of PSNR, SSIM, and LPIPS. denotes numbers taken from [137].
[…]
Mip-NeRF [268] [ICCV21] 18.08 0.57 0.55
Mip-NeRF 360 [8] [CVPR22] 22.61 0.69 0.40
Instant-NGP [9] [TOG22] 16.78 0.52 0.57
Urban-NeRF [269] [CVPR22] 20.75 0.63 0.48
S-NeRF [270] [ICLR23] 25.43 0.73 0.30
SUDS [271] [CVPR23] 21.26 0.60 0.47
3D GS [10] [TOG23] 26.08 0.72 0.30
DrivingGaussian-L [137] [CVPR24] 28.74 0.86 0.24
6.4 Performance Benchmarking: Driving Scenes
This section focuses on evaluating the rendering quality in driving scenes, which is essential for autonomous driving.
$\bullet$ Dataset: nuScenes [243] dataset is a comprehensive collection for autonomous driving, featuring 1000 driving scenes captured using an array of sensors including six cameras, one LiDAR, five RADARs, GPS, and IMU. It provides detailed annotations for 23 object classes with 3D bounding boxes. Six challenging scenes are used for evaluation [137].
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve one 3D GS based approach [137], as well as six NeRF based methods [268, 8, 9, 269, 270, 271].
$\bullet$ Evaluation Metric: PSNR, SSIM [263], and LPIPS [264] are used for evaluation.
$\bullet$ Result: The results in Table IV demonstrate that the 3D GS based methods significantly surpass the NeRF based methods across all evaluated metrics. For instance, DrivingGaussian-L [137] outperforms S-NeRF [270] by 3.31dB in terms of PSNR. This suggests that 3D Gaussians can benefit from multi-sensor information to capture dynamic objects in driving scenes, especially fast-moving ones.
@@ 6.5 Performance Benchmarking: Human Avatar
$\bullet$ Dataset: ZJU-MoCap [272] dataset is a prevalent benchmark in human modeling from videos, captured with 23 synchronized cameras at a 1024$\times$1024 resolution. Following [273], six subjects (i.e., 377, 386, 387, 392, 393, and 394) are used for evaluation.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent papers which model human avatar with 3D GS [210, 224, 212, 221], as well as six human rendering approaches [272, 274, 275, 276, 273, 277].
@@
TABLE V: Quantitative avatar modeling results6.5) on ZJU-MoCap [272], in terms of PSNR, SSIM, and LPIPS*. The numbers of non-GS methods and FPS are taken from [210] and [224], respectively. denotes the average of the values reported in the original paper.
[…]
NeuralBody [272] [CVPR21] 29.03 0.96 42.47 3.5
AnimNeRF [274] [ICCV21] 29.77 0.96 46.89 2.1
PixelNeRF [275] [ICCV21] 24.71 0.89 121.86 -
NHP [276] [NeurIPS21] 28.25 0.95 64.77 -
HumanNeRF [273] [CVPR22] 30.66 0.97 33.38 0.4
Instant-NVR [277] [CVPR23] 31.01 0.97 38.45 1.5
Human101 [224] [arXiv] 31.79 0.96 35.75 104.0
HUGS [212] [CVPR24] 30.98 0.97 26.67 -
3DGS-Avatar [221] [CVPR24] 30.61 0.97 29.58 -
GART [210] [CVPR24] 32.22 0.98 29.21 -
$\bullet$ Result: Table V presents the numerical results of top-leading solutions in human avatar modeling. We observe that introducing 3D GS into the framework leads to consistent performance improvements in both rendering quality and speed. For instance, GART [210] outperforms current SOTA, Instant-NVR [277], by 1.21dB in terms of PSNR. Note that Human101 [224] is about 68 times faster in achieving very competitive accuracy compared to Instant-NVR [277]. Considering the enhanced fidelity, inference speed and editability, 3D GS based avatar modeling may revolutionize the field of 3D animation, interactive gaming, etc.
@@ 6.6 Performance Benchmarking: Surgical Scenes
$\bullet$ Dataset: EndoNeRF [252] dataset presents a specialized collection of stereo camera captures, comprising two samples of in-vivo prostatectomy. It is tailored to represent real-world surgical complexities, including challenging scenes with tool occlusion and pronounced non-rigid deformation.
@@
TABLE VI: Quantitative surgical 3D reconstruction results6.6) on EndoNeRF [252], in terms of PSNR, SSIM, and LPIPS. The numbers of non-GS methods, FPS, and GPU usage (Mem.) are taken from [141]. denotes numbers taken from [137]. denotes the average of the values reported in the original paper.
@@ 7 Future Research Directions
$\bullet$ Physics- and Semantics-aware Scene Representation. As a new, explicit scene representation technique, 3D Gaussian offers transformative potential beyond merely enhancing novel-view synthesis. It has the potential to pave the way for simultaneous advancements in scene reconstruction and understanding by devising physics- and semantics-aware 3D GS systems. This is poised to revolutionize a range of fields and downstream applications. For instance, incorporating prior knowledge such as the general shape of objects can reduce the need for extensive training viewpoints [48, 47] while improving geometry/surface reconstruction [100, 77]. A critical metric for assessing scene representation is the realism of its generated scenes, which encompasses challenges in geometry, texture, and lighting fidelity [121, 134, 66]. By merging physical principles and semantic information within the 3D GS framework, one can expect that the realism will be enhanced, thereby facilitating dynamics modeling [278, 21], editing [92, 90], generation [126, 127], and beyond. In a nutshell, pursuing this advanced and versatile scene representation opens up new possibilities for innovation in computational creativity and practical applications across diverse domains.
@@
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these physical priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and realism of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the Spring-Mass 3D Gaussians for elastic object modeling [279], and the generalizable Gaussian representation based on multi-view stereo [280]. Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
$\bullet$ 3D GS for Robotics. In the world of robotics, especially when it comes to robots performing tasks that involve handling objects in ways similar to humans, there is a growing need for these machines to be able to navigate and manipulate their environment in a more intuitive and dynamic manner. This necessity stems from the desire to deploy intelligent robots in real-world settings, where they are often faced with new and unfamiliar tasks. Traditional approaches to robotic manipulation have relied heavily on understanding the environment through semantic representation, which means identifying objects and their attributes. However, these methods often overlook the importance of how things move and interact over time, which is crucial for completing tasks in the way humans intend. Imagine a robot trying to stack blocks based on a verbal command. Using only semantic representation, the robot might recognize the blocks but fail to stack them if it does not understand how the blocks should be positioned and moved relative to each other over time. This gap in understanding can lead to failure in executing tasks that require interaction with multiple objects or navigating dynamic environments. Due to its explicit nature, 3D GS can be used beyond mere semantic and structural analysis of environments; it also encompasses the dynamic aspect, providing a comprehensive understanding of how scenes evolve and objects interact over time. This is essential for robots tasked with navigating and manipulating their surroundings. Although initial GS based efforts on world models [281, 282] and reinforcement learning [283, 284] are promising, they represent only the beginning of what is possible. Further research in this area is anticipated to enhance the capabilities of robots in performing tasks that require an understanding of both the physical space and the temporal changes within it.
@@
$\bullet$ Modeling Internal Structures of Objects with 3D GS. Despite the ability of 3D GS to produce highly realistic renderings, modeling internal structures of objects (e.g., for a scanned object in computed tomography) within the current GS framework presents a notable challenge. Due to the splatting and density control process, the current representation of 3D Gaussian is unorganized and cannot align well with the object’s actual internal structures. Moreover, there is a strong preference in various applications to depict objects as volumes (e.g., computed tomography). However, the disordered nature of 3D GS makes volume modeling particularly difficult. Li et al. [285] used 3D Gaussians with density control as the basis for the volumetric representation and did not involve the splatting process. X-Gaussian [286] involves the splatting process for fast training and inference but cannot generate volumetric representation. Using 3D GS to model the internal structures of objects remains unanswered and deserves further exploration.
@@
$\bullet$ 3D GS for Simulation in Autonomous Driving. Collecting real-world datasets for autonomous driving is both expensive and logistically challenging, yet crucial for training effective image-based perception systems. To mitigate these issues, simulation emerges as a cost-effective alternative, enabling the generation of synthetic datasets across diverse environments. However, the development of simulators capable of producing realistic and diverse synthetic data is fraught with challenges. These include achieving a high level of realism, accommodating various control methods, and accurately simulating a range of lighting conditions. While early efforts [137, 138, 139] in reconstructing urban/street scenes with 3D GS have been encouraging, they are just the tip of the iceberg in terms of the full capabilities. There remain numerous critical aspects to be explored, such as the integration of user-defined object models, the modeling of physics-aware scene changes (e.g., the rotation of vehicle wheels), and the enhancement of controllability and realism (e.g., in varying lighting conditions).

### v5

@@ 6 Performance Comparison
In this section, we provide more empirical evidence by presenting the performance of several 3D GS algorithms that we previously discussed. The diverse applications of 3D GS across numerous tasks, coupled with the custom-tailored algorithmic designs for each task, render a uniform comparison of all 3D GS algorithms across a single task or dataset impracticable. For comprehensiveness, we provide a collection of representative datasets in Table II according to our analysis in Sec. 5. Due to the limited space, we have chosen several representative tasks for an in-depth performance evaluation. The performance scores are primarily sourced from the original papers, except where indicated otherwise. We also maintain a Github repository for this section.
@@ 6.1 Performance Benchmarking: Localization
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve four recent 3D GS based algorithms [114, 113, 112, 111] and six typical SLAM methods [262, 263, 264, 265, 266, 267].
@@
$\bullet$ Result: As shown in Table I, the recent 3D Gaussians based localization algorithms have a clear advantage over existing NeRF based dense visual SLAM. For example, SplaTAM [112] achieves a trajectory error improvement of $\sim$50%, decreasing it from 0.52cm to 0.36cm compared to the previous state-of-the-art (SOTA) [266]. We attribute this to the dense and accurate 3D Gaussians reconstructed for scenes, which can handle the noise of real sensors. This reveals that effective scene representations can improve the accuracy of localization tasks.
@@ 6.2 Performance Benchmarking: Static Scenes
$\bullet$ Dataset: The same dataset as in Sec. 6.1, i.e., Replica [261], is used for comparison. The testing views are the same as those collected by [262].
@@
$\bullet$ Result: Table III shows that 3D Gaussians based systems generally outperform the three dense SLAM competitors. For example, Gaussian-SLAM [114] establishes new SOTA and outperforms previous methods by a large margin. Compared to Point-SLAM [266], GSSLAM [113] is about 578 times faster in achieving very competitive accuracy. In contrast to previous method [266] that relies on depth information, such as depth-guided ray sampling, for synthesizing novel views, 3D GS based system eliminates this need, allowing for high-fidelity rendering for any views.
@@ 6.3 Performance Benchmarking: Dynamic Scenes
$\bullet$ Dataset: D-NeRF [184] dataset includes videos with 50 to 200 frames each, captured from unique viewpoints. It features synthetic, animated objects in complex scenes, with non-Lambertian materials. The dataset provides 50 to 200 training images and 20 test images per scene, designed for evaluating models in the monocular setting. The testing views are the same as the original paper [184].
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve five recent papers that model dynamic scenes with 3D GS [126, 95, 132, 94, 93], as well as six NeRF based approaches [184, 302, 37, 303, 187, 304].
@@
$\bullet$ Result: From Table IV we can observe that 3D GS based methods outperform existing SOTAs by a clear margin. The static version of 3D GS [10] fails to reconstruct dynamic scenes, resulting in a sharp drop in performance. By modeling the dynamics, D-3DGS [94] outperforms the SOTA method, FFDNeRF [187], by 6.83dB in terms of PSNR. These results indicate the effectiveness of introducing additional properties or structured information to model the deformation of Gaussians so as to model the scene dynamics.
TABLE IV: Comparison of reconstruction methods (§6.3) on D-NeRF [184] (dynamic scenes), in terms of PSNR, SSIM, and LPIPS. denotes results reported in [95].
@@ 6.4 Performance Benchmarking: Human Avatar
$\bullet$ Dataset: ZJU-MoCap [292] is a prevalent benchmark in human modeling from videos, captured with 23 synchronized cameras at a 1024$\times$1024 resolution. Six subjects (i.e., 377, 386, 387, 392, 393, and 394) are used for evaluation [305]. The same testing views following [306] are adopted.
$\bullet$ Benchmarking Algorithms: For performance comparison, we involve three recent papers which model human avatar with 3D GS [146, 145, 249], as well as six human rendering approaches [292, 307, 308, 309, 305, 306].
@@
$\bullet$ Result: Table V presents the numerical results of top-leading solutions in human avatar modeling. We observe that introducing 3D GS into the framework leads to consistent performance improvements in both rendering quality and speed. For instance, GART [146] outperforms current SOTA, Instant-NVR [306], by 1.21dB in terms of PSNR. Considering the enhanced fidelity, inference speed and editability, 3D GS based avatar modeling may revolutionize the field of 3D animation, interactive gaming, etc.
@@ 6.5 Performance Benchmarking: Surgical Scenes
$\bullet$ Dataset: EndoNeRF [258] dataset presents a specialized collection of stereo camera captures, comprising two samples of in-vivo prostatectomy. It is tailored to represent real-world surgical complexities, including challenging scenes with tool occlusion and pronounced non-rigid deformation. The same testing views as in [260] are used.
@@
TABLE V: Comparison of reconstruction methods (§6.4) on ZJU-MoCap [292] (avatar), in terms of PSNR, SSIM, and LPIPS*. The results for non-GS methods are taken from [146].
[…]
NeuralBody [292] [CVPR21] 29.03 0.96 42.47
AnimNeRF [307] [ICCV21] 29.77 0.96 46.89
PixelNeRF [308] [ICCV21] 24.71 0.89 121.86
NHP [309] [NeurIPS21] 28.25 0.95 64.77
HumanNeRF [305] [CVPR22] 30.66 0.97 33.38
Instant-NVR [306] [CVPR23] 31.01 0.97 38.45
GauHuman [145] [CVPR24] 31.34 0.97 30.51
3DGS-Avatar [249] [CVPR24] 30.61 0.97 29.58
GART [146] [CVPR24] 32.22 0.98 29.21
TABLE VI: Comparison of reconstruction methods (§6.5) on EndoNeRF [258] (surgical scenes), in terms of PSNR, SSIM, and LPIPS. The results for non-GS methods are taken from [153]. FPS and GPU usage for training (Mem.) are measured using one 4090 GPU [153].
@@ 7 Future Research Directions
$\bullet$ Physics- and Semantics-aware Scene Representation. As a new, explicit scene representation technique, 3D Gaussian offers transformative potential beyond merely enhancing novel-view synthesis. It has the potential to pave the way for simultaneous advancements in scene reconstruction and understanding by devising physics- and semantics-aware 3D GS systems. While significant progress has been made in physics (Sec. 5.7) and semantics [310, 311, 312, 313, 314, 315] individually, there remains considerable untapped potential in their synergistic integration. This is poised to revolutionize a range of fields and downstream applications. For instance, incorporating prior knowledge such as the general shape of objects can reduce the need for extensive training viewpoints [48, 47] while improving geometry/surface reconstruction [316, 77]. A critical metric for assessing scene representation is the quality of its generated scenes, which encompasses challenges in geometry, texture, and lighting fidelity [128, 141, 66]. By merging physical principles and semantic information within the 3D GS framework, one can expect that the quality will be enhanced, thereby facilitating dynamics modeling [166, 21], editing [92, 90], generation [133, 134], and beyond. In a nutshell, pursuing this advanced and versatile scene representation opens up new possibilities for innovation in computational creativity and practical applications across diverse domains.
@@
$\bullet$ Learning Physical Priors from Large-scale Data. As we explore the potential of physics- and semantics-aware scene representations, leveraging large-scale datasets to learn generalizable, physical priors emerges as a promising direction. The goal is to model the inherent physical properties and dynamics embedded within real-world data, transforming them into actionable insights that can be applied across various domains such as robotics and visual effects. Establishing a learning framework for extracting these generalizable priors enables the application of these insights to specific tasks in a few-shot manner. For instance, it allows for rapid adaptation to new objects and environments with minimal data input. Furthermore, integrating physical priors can enhance not only the accuracy and quality of generated scenes but also their interactive and dynamic qualities. This is particularly valuable in AR/VR environments, where users interact with virtual objects that behave in ways consistent with their real-world counterparts. However, the existing body of work on capturing and distilling physics-based knowledge from extensive 2D and 3D datasets remains sparse. Notable efforts in related area include the continuum mechanics based GS systems (Sec. 5.7), and the generalizable Gaussian representation based on multi-view stereo [317]. Further exploration on real2sim and sim2real might offer viable routes for advancements in this field.
@@
$\bullet$ Modeling Internal Structures of Objects with 3D GS. Despite the ability of 3D GS to produce highly photorealistic renderings, modeling internal structures of objects (e.g., for a scanned object in computed tomography) within the current GS framework presents a notable challenge. Due to the splatting and density control process, the current representation of 3D Gaussian is unorganized and cannot align well with the object’s actual internal structures. Moreover, there is a strong preference in various applications to depict objects as volumes (e.g., computed tomography). However, the disordered nature of 3D GS makes volume modeling particularly difficult. Li et al. [318] used 3D Gaussians with density control as the basis for the volumetric representation and did not involve the splatting process. X-Gaussian [319] involves the splatting process for fast training and inference but cannot generate volumetric representation. Using 3D GS to model the internal structures of objects remains unanswered and deserves further exploration.
@@
$\bullet$ 3D GS for Simulation in Autonomous Driving and beyond. Collecting real-world datasets for autonomous driving is both expensive and logistically challenging, yet crucial for training effective image-based perception systems. To mitigate these issues, simulation emerges as a cost-effective alternative, enabling the generation of synthetic datasets across diverse environments. However, the development of simulators capable of producing photorealistic and diverse synthetic data is fraught with challenges. These include achieving a high level of quality, accommodating various control methods, and accurately simulating a range of lighting conditions. While early efforts [188, 189, 190] in reconstructing urban/street scenes with 3D GS have been encouraging, they are just the tip of the iceberg in terms of the full capabilities. There remain numerous critical aspects to be explored, such as the integration of user-defined object models, the modeling of physics-aware scene changes (e.g., the rotation of vehicle wheels), and the enhancement of controllability and quality (e.g., in varying lighting conditions). Mastery of these capabilities would not only advance autonomous systems but also redefine computational understanding of physical spaces — a leap with implications for world models, spatial intelligence, embodied AI, and beyond.